OpenStudy (anonymous):

Simple interest on a certain sum of money for 4 years at 4% per annum exceeds the compound interest on the same sum for 3 years at 5% per annum by Rs. 228. Find the sum.

OpenStudy (anonymous):

16x/100-x(105/100)3-x=228

OpenStudy (anonymous):

x=12000

ganeshie8 (ganeshie8):

simple interest for 4 years : \(\large x*\dfrac{4}{100}*4\) compound interest for 3 years : \(\large x(1+0.05)^3-x \)

ganeshie8 (ganeshie8):

difference : \(\large \left( x*0.04*4\right) - \left( x(1+0.05)^3-x\right) \gt 228\)

ganeshie8 (ganeshie8):

you should get x = 96,000 check ur calculation again